OpenStudy (anonymous):

how do I factor 8x^2+21=-59x

OpenStudy (anonymous):

First add 59x to the other side so that 8x^2 + 59x + 21 = 0 Then find factors that multiply to 168 and add up to 59. Do you know how to do the rest?

OpenStudy (anonymous):

The factors are 56 and 3. Reinsert them in the place of 59: 8x^2 + 56x+ 3x + 21 = 0. Now group the first 2 terms together and group the second 2 terms together by placing parentheses around them. (8x^2 + 56x) + (3x + 21) = 0 Now factor out the GCF separately in both binomials. 8x(x + 7) + 3(x + 7). Now combine the numbers outside the parentheses and the numbers inside separately. The final answer is (8x + 3)(x + 7)